Bridgeport Co., a manufacturer of rain barrels, had the following data for 2016.
Sales 3,300 units
Sales price $100 per unit
Variable costs $37 per unit
Fixed costs $172,557

Collapse question part

(a)

What is the contribution margin ratio? (Round answer to 0 decimal places, e.g. 5,275.)
Contribution margin ratio

%
Click if you would like to Show Work for this question:
Answer #1

Contribution margin=Sales-Variable cost

=(100-37)=$63 per unit

Contribution margin ratio=Contribution margin/Sales

=63/100

=63%